# A New Benchmark and Model for Challenging Image Manipulation Detection

### v1

@@ Conclusion
This study presents a novel Challenging Image Manipulation Detection (CIMD) dataset, which comprise of two subsets that are designed for evaluating image-editing-based and compression-based approaches, respectively. The datasets were manually taken and tampered with, and come with high-quality annotations. Additionally, we propose a two-branch method that outperforms state-of-the-art models in detecting image manipulations using the CIMD dataset. We will release our dataset to facilitate future research.
Future Work includes the exploration of the identification of large tampering regions ($\geqslant$ 90%).

### v2

@@ Conclusion
This study presents a novel Challenging Image Manipulation Detection (CIMD) dataset, which comprise of two subsets that are designed for evaluating image-editing-based and compression-based approaches, respectively. The datasets were manually taken and tampered with, and come with high-quality annotations. Additionally, we propose a two-branch method that outperforms state-of-the-art models in detecting image manipulations using the CIMD dataset. We have released our dataset to facilitate future research.
Ethics Statement
To ensure ethical compliance, all photos presented in our dataset are original and obtained either in public places or with the owners’ explicit permission in private places, in accordance with local jurisdiction laws. Moreover, the authors ensure that the photos contain neither identifiable individuals nor personal information. As advised by institutional review boards (IRB), IRB approval is not required for the dataset.